### v3

@@ 5.2 Significance of Bright
In this paper, we introduced Bright, the first globally distributed, multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. Bright includes pre-event optical imagery and post-event SAR imagery with sub-meter spatial resolution. Beyond introducing the dataset, we conducted a comprehensive series of experiments to validate its utility. We benchmarked several state-of-the-art supervised learning models under a standard machine learning split, and further extended the evaluation to a cross-event transfer setting, simulating real-world scenarios where no/limited target annotations are available. Additionally, we assessed the performance of unsupervised domain adaptation and semi-supervised learning methods, as well as unsupervised multimodal change detection and image matching techniques. These findings provide both performance baselines and valuable insights into model design for future research. Bright is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to further improve Bright’s utility for future research within the community and for practical disaster response applications at regional, national, and international levels.
[…]
The Bright dataset and the code for training and testing benchmark methods (including code related to IEEE GRSS DFC 2025) are available at https://github.com/ChenHongruixuan/BRIGHT.
Appendix A Feature visualization
@@
For the feature visualisation of buildings in different events in Figure 5c, we specifically employ DINOv2 (Oquab et al., 2024) to extract the high-dimensional features of buildings in different events. The corresponding reference maps are used to exclude background features. The features are then visualized using the t-SNE algorithm (van der Maaten and Hinton, 2008).
@@
Appendix B Manual registration and estimating registration errors
Figure 11: Overview of manually selected control points used for registration on some events. The number in parentheses of each event is the number of control points selected on that view of the image. The labeled yellow boxes highlight specific zoom-in regions that are shown in below.
The manually selected control points by EO experts on some disaster scenes are shown in Figure 11. Despite multiple rounds of meticulous registration and cross-validation by several EO experts, registration errors between optical and SAR images cannot be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we propose a proxy method for registration error estimation:
@@ Appendix B Manual registration and estimating registration errors
Appendix C Details of benchmark DL models
@@
  • For UNet, the encoder has five convolutional blocks, each consisting of two 3$\times$3 convolutional layers and two batch normalization layers. The number of channels in the five blocks are 64, 128, 256, 512, and 1024, respectively. The decoder contains four convolutional blocks, and the number of channels is set accordingly to the number of channels of the skip-connected features extracted by the encoder.
@@
  • For SiamCRNN, we applied pseudo-siamese ResNet-18 as encoder and four ConvLSTM layers (Shi et al., 2015) with 3$\times$3 convolutional kernel size and a hidden dimension of 128, as the decoder.
@@
Appendix D Details of post-processing methods
To further refine the raw predictions from deep learning models and enhance the quality of building damage maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 4.3. This section provides an overview of these methods and details the specific implementation used in our study.
@@ Appendix D Details of post-processing methods
  • Object-based major voting aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
@@
Appendix E Details of unsupervised domain adaptation methods
@@
Appendix F Details of semi-supervised learning methods
@@
Final performance is evaluated on the test set of the target domain using mIoU as the primary evaluation metric.
Appendix G Details of unsupervised multimodal change detection methods and a more practical evaluation protocol
Figure 12: Illustration of benchmark datasets commonly used in unsupervised multimodal change detection. The order from left to right is the corresponding order in the Table 11. In reference maps, white color indicates changed pixels, and black color indicates unchanged pixels.
@@ Appendix G Details of unsupervised multimodal change detection methods and a more practical evaluation protocol
The training and evaluation protocols commonly adopted in the UMCD literature are often limited to individual image pairs (shown in Figure 12), using the same scene for representation learning, hyperparameter tuning, and evaluation. While this is understandable due to the lack of large-scale public datasets in this field, such a setup fails to reflect real-world use cases and often leads to overfitting and overestimated performance. To address this, we introduce a standardized and practical evaluation protocol using the Bright dataset. Specifically, we use the validation set from the standard ML split in Section 2.3 as the training set for UMCD methods, including any hyperparameter or threshold tuning. These models are then evaluated on the combined training and test sets, which are strictly held out during adaptation. This avoids overlap between tuning and evaluation scenes and better simulates the real-world deployment setting, where a model is expected to generalize to new, unseen disasters.
For evaluation, it is important to note that most of these methods were originally developed for land cover change detection, where changes correspond to transitions between semantic categories such as vegetation, water, or built-up areas. Applying them directly to disaster damage detection poses certain challenges. In particular, “Damaged” buildings do not always exhibit strong spatial or spectral signals, making them hard to distinguish in an unsupervised setting. However, “Destroyed” buildings often result in a complete change of land cover appearance (e.g., collapsed structures, debris, or ground exposure), which aligns better with the assumptions of existing UMCD methods. Therefore, in our evaluation, we treat buildings labeled as “Destroyed” in Bright as the positive (changed) class, and all other regions (including intact, damaged, and background) as the unchanged class. This allows a fair adaptation of these methods to a disaster-relevant setting while respecting their original design.
Appendix H Details of unsupervised multimodal image matching methods
@@
In contrast, area-based methods, including FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019), begin by identifying keypoints only in the optical image. These points are then matched by sliding a template window across a local search region in the SAR image, relying on local patch similarity. In our experiments, we used a template size of 120$\times$120 pixels, and a search window of 200$\times$200 pixels for both methods. This category is also implemented using open-source code [^11] https://github.com/yeyuanxin110/CFOG.
@@
where $(x_{i}^{\text{sar}},y_{i}^{\text{sar}})=\mathcal{T}(x_{i}^{\text{opt}},y_{i}^{\text{opt}})$ are transformed SAR coordinates.
@@
Hongruixuan Chen: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (equal). Jian Song: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). Olivier Dietrich: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). Clifford Broni-Bediako: conceptualization (support), data curation (support), writing – review and editing (lead). Weihao Xuan: data curation (support), project administration (support), writing – review and editing (equal). Junjue Wang: data curation (support), writing – review and editing (support). Xinlei Shao: funding acquisition (support), writing – review and editing (support). Yimin Wei: writing – review and editing (support). Junshi Xia: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). Cuiling Lan [^12] Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). Konrad Schindler: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). Naoto Yokoya: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).

### v4

@@ 5 Data and Code Availability
In this paper, we introduced Bright, the first globally distributed multimodal dataset with open access to the community, covering 14 natural and human-made disaster events. Bright includes pre-event optical and post-event SAR images with sub-meter spatial resolution. Beyond introducing the dataset, we conducted a comprehensive series of experiments to validate its utility. We benchmarked several state-of-the-art supervised learning models under a standard machine learning data split. Moreover, we extended the evaluation to a cross-event transfer setting, simulating real-world scenarios where no or limited target annotations are available. Furthermore, we assessed the performance of unsupervised domain adaptation, semi-supervised learning methods, unsupervised multimodal change detection, and image matching techniques. The findings serve as performance baselines and provide valuable insights for future research in DL model design for real-world disaster response. Bright is an ongoing project, and we remain committed to continuously enhancing its diversity and quality by incorporating new disaster events and refining the existing data. Our objective is to improve Bright’s utility for practical disaster response applications at all levels (regional, national, and international) and research in the community.
Appendix A Details of disaster events
A.1 Explosion in Beirut, 2020
On August 4, 2020, a massive explosion occurred at the Port of Beirut in Lebanon, caused by the improper storage of 2,750 tons of ammonium nitrate. The explosion caused widespread damage within a radius of several kilometers, significantly impacting the port and surrounding neighborhoods, including areas such as Gemmayzeh, Mar Mikhael, and Achrafieh. It resulted in 218 deaths and more than 7,000 injuries, and left approximately 300,000 people homeless (Fakih and Majzoub, 2021). Economic losses were estimated to be 15 billion USD (Valsamos et al., 2021). The disaster compounded Lebanon’s ongoing economic challenges and contributed to political instability and social unrest.
A.2 Explosion in Bata, 2021
On March 7, 2021, a series of four explosions occurred at the Cuartel Militar de Nkoantoma in Bata, Equatorial Guinea, caused by improperly stored explosives. The blasts led to at least 107 deaths and over 615 injuries, and widespread destruction throughout the city (OCHA, 2021a). A total of 243 structures were destroyed or severely damaged, displacing many residents. Around 150 families sought refuge in temporary shelters, while others stayed with relatives. Local hospitals treated more than 500 injured individuals, and the economic impact was severe, underscoring the dangers associated with the improper storage of hazardous materials.
A.3 Volcano Eruption in DR Congo and Rwanda, 2021
On May 22, 2021, Mount Nyiragongo in the Democratic Republic of the Congo erupted, causing widespread devastation. The eruption resulted in 32 deaths and the destruction of 1,000 homes. The displacement of thousands as lava flows threatened the city of Goma (IFRC, 2021). Nearly 400,000 people were evacuated due to the risk of further volcanic activity, including potential magma flow beneath Goma and nearby Lake Kivu. Despite continued seismic activity, life in Goma largely returned to normal by August 2021. However, plans to relocate parts of the city remain under consideration due to the ongoing threat from the volcano.
A.4 Earthquake in Haiti, 2021
On August 14, 2021, a magnitude 7.2 earthquake struck Haiti’s Tiburon Peninsula, primarily affecting the Nippes, Sud, and Grand’Anse departments. The disaster caused over 2,200 deaths and more than 12,200 injuries, and left thousands homeless. The economic losses were significant, estimated at over USD 1.5 billion. Approximately 137,500 buildings, including homes, schools, and hospitals, were damaged or destroyed (OCHA, 2021b). As the deadliest natural disaster of 2021, the earthquake exacerbated Haiti’s existing challenges, including widespread poverty and political instability.
A.5 Volcano Eruption in La Palma, 2021
On September 19, 2021, the Cumbre Vieja volcano on La Palma, part of Spain’s Canary Islands, erupted following several days of seismic activity. The eruption primarily impacted the island’s western side, covering over 1,000 hectares with lava and destroying more than 3,000 buildings, including the towns of Todoque and La Laguna. The lava flow, measuring about 3.5 kilometers wide and 6.2 kilometers long, reached the sea, cutting off the coastal highway and forming a new peninsula with extensive lava tubes. Although the timely evacuation of around 8,000 people prevented major casualties, one person died from inhaling toxic gases (Troll et al., 2024). It caused significant damage to arable land and affected livelihoods, displacing thousands of residents. Economic losses exceeded 800 million Euros.
A.6 Wildfire in Colorado, 2021
The Marshall Fire, which started on December 30, 2021, in Boulder County, Colorado, became the most destructive wildfire in the State’s history in terms of destroyed buildings. Fueled by dry grass from an unusually warm and dry season, and winds up to 185 km/h, the fire killed two people and injured eight. It destroyed 1,084 structures, including homes, a hotel, and a shopping center, causing over 2 billion USD in damage (Forrister et al., 2024). More than 37,500 residents were evacuated, and significant damage to public drinking water systems occurred (Forrister et al., 2024).
A.7 Armed Conflict in Ukraine, 2022
In February 2022, Russian forces launched a full-scale invasion of Ukraine, resulting in widespread destruction and displacement. By November 2024, total damages to Ukraine’s infrastructure had reached \$170 billion. More than 236,000 residential buildings were damaged or destroyed, including 209,000 private houses, 27,000 apartment buildings, and 600 dormitories. Over 4,000 educational institutions and 1,554 medical facilities were also affected, with extensive damage to transport, energy, and telecommunications infrastructure. The conflict has inflicted more than 40,000 civilian casualties, displaced four million people internally, and forced 6.8 millionto flee. As of late 2024, approximately 14.6 million Ukrainians require humanitarian assistance (Andrienko et al., 2025).
A.8 Earthquake in Turkey, 2023
On February 6, 2023, a magnitude 7.8 earthquake struck southeastern Turkey near Gaziantep, followed by a magnitude 7.7 aftershock. The disaster, the most powerful earthquake in Turkey since 1939, caused widespread destruction across approximately 350,000 $km^{2}$, affecting 14 million people and displacing 1.5 million. The death toll reached 53,537, with 107,213 injuries (STL, 2024), making it one of the deadliest earthquakes in modern history. Economic losses were estimated at 148.8 billion USD, with over 518,000 houses and 345,000 apartments destroyed (Government of Türkiye, 2023). The earthquake caused severe damage to infrastructure, agriculture, and essential services, further worsening the region’s economic challenges. International aid was mobilized to support the affected populations.
A.9 Cyclone in Myanmar, 2023
In May 2023, Cyclone Mocha, a Category 5 hurricane, struck Myanmar, causing widespread devastation in the country’s coastal regions, particularly in Rakhine State. According to official estimates, at least 148 people were killed and 132 injured, although other sources suggest higher figures (GFDRR and World Bank, 2023). The cyclone affected around 1.2 million people in Rakhine alone, with over 200,000 buildings reportedly damaged or destroyed, making it one of the most destructive cyclones in the region in the past 15 years. The direct economic damage was estimated at 2.24 billion USD in damages, equivalent to 3.4% of Myanmar’s GDP (GFDRR and World Bank, 2023).
A.10 Wildfire in Hawaii, 2023
In August 2023, a series of wildfires broke out on the island of Maui, Hawaii, causing widespread destruction and significant impacts on the local population and environment. The fires, fueled by dry conditions and strong winds, primarily affected the town of Lahaina, where at least 102 people were killed and two remain missing (Hedayati et al., 2024). Over 2,200 buildings were destroyed, including many historic landmarks, resulting in estimated damages of 5.5 billion USD (Jones et al., 2024; NCEI, 2025). The fires prompted evacuations and led to the displacement of thousands of residents, with significant economic losses in the tourism and agriculture sectors.
A.11 Earthquake in Morocco, 2023
On September 8, 2023, a 6.9 magnitude earthquake struck Morocco’s Al Haouz Province near Marrakesh, causing widespread devastation. The earthquake killed nearly 3,000 people, injured more than 5,500, and displaced more than half a million (OCHA, 2023). The earthquake damaged or destroyed nearly 60,000 houses, with the heaviest losses reported in rural communities of the Atlas Mountains (International Federation of Red Cross and Red Crescent Societies, 2024). Overall damage is estimated at about USD 7 billion (National Geophysical Data Center / World Data Service, 2025), while direct economic losses of roughly USD 30 million amount to approximately $0.24\%$ of Morocco’s GDP. Cultural heritage in Marrakesh was also hard‑hit, sections of the UNESCO‑listed Medina and several historic mosques sustained severe damage.
A.12 Flood and Storm in Libya, 2023
In September 2023, Storm Daniel brought catastrophic flooding to northeastern Libya, particularly in Derna, when two dams collapsed and released an estimated 30 million cubic meters of water. At least 5,923 people were killed, though local officials warned fatalities could reach 18,000–20,000 (USAID Bureau for Humanitarian Assistance, 2024). Post–event analyses revealed that approximately $10\%$ of Derna’s housing stock was destroyed and a further $18.5\%$ sustained damage. Across the wider coastal belt—from Benghazi, through Jabal Al Akhdar and Al Marj, to Derna—an estimated 18,838 dwellings were damaged or obliterated (Normand and Heggy, 2024). The disaster, considered the second deadliest dam failure in history, destroyed the city’s infrastructure; four bridges collapsed, and entire neighborhoods were washed out to sea. Long-standing neglect of dam maintenance, compounded by Libya’s political turmoil, contributed significantly to the scale of the tragedy.
A.13 Hurricane in Mexico, 2023
In September 2023, Hurricane Norma, a Category 4 hurricane, struck the western coast of Mexico, severely affecting Sinaloa and Baja California Sur. This was followed by Hurricane Otis in October, which made landfall near Acapulco as a Category 5 hurricane. Otis was the strongest Pacific hurricane to hit Mexico, causing at least 52 deaths and leaving 32 missing. The storm caused unprecedented destruction, with more than 51,864 homes destroyed and damages estimated at 12-16 billion USD, surpassing Hurricane Wilma as the costliest Mexican hurricane (Reinhart and Reinhart, 2024).
A.14 Earthquake in Noto, 2024
On January 1, 2024, a magnitude 7.5 earthquake struck the Noto Peninsula in Ishikawa Prefecture, Japan, reaching a maximum JMA seismic intensity of Shindo 7. The earthquake caused widespread destruction, particularly in the towns of Suzu, Wajima, Noto, and Anamizu, and triggered a 7.45-meter tsunami along the coast of the Sea of Japan. The disaster resulted in 572 deaths and over 1,300 injuries, with 193,529 structures damaged across nine prefectures (Fire and Disaster Management Agency, 2025). It was the deadliest earthquake in Japan since the 2011 Tōhoku disaster. As of February 20 2024, 12,929 people remained in 521 evacuation centers (Japanese Red Cross Society, 2024). The event prompted Japan’s first major tsunami warning since 2011. Total damage is estimated at 16.9 trillion Japanese Yen, reflecting substantial economic losses with significant impacts on infrastructure and communities across the affected regions (Kaneko, 2024).
Appendix B Feature visualization
@@
For the feature visualisation of buildings in different events in Figure 5-(c), we specifically employ DINOv2 (Oquab et al., 2024) to extract the high-dimensional features of buildings in different events. The corresponding reference maps are used to exclude background features. The features are then visualized using the t-SNE algorithm (van der Maaten and Hinton, 2008).
@@
Appendix C Manual registration and estimating registration errors
Figure 12: Overview of manually selected control points used for registration on some events. The number in parentheses of each event is the number of control points selected on that view of the image. The labeled yellow boxes highlight specific zoom-in regions that are shown in below. Data source: SAR images of Bata-Explosion-2020 and La Palma-Volcano-2020 are sourced from Capella Space Open Data Gallery while the SAR image of Noto-Earthquake-2024 is from Umbra Open Data Program.
We performed the manual registration process using QGIS, with the “Georeferencer” plugin to align SAR images to the optical imagery as the reference. The transformation type was set to “Thin Plate Spline”, and “Lanczos resampling (6×6 kernels)” was applied to achieve high-quality interpolation. The manually selected control points by EO experts on some disaster scenes are shown in Figure 12.
Despite multiple rounds of meticulous registration and cross-validation by several EO experts, registration errors between optical and SAR images cannot be completely eliminated. Since we cannot directly obtain ground truth (i.e., actual ground control points), we propose a proxy method for registration error estimation:
@@ Appendix C Manual registration and estimating registration errors
Appendix D Details of benchmark deep learning models
@@
  • For UNet, the encoder has five convolutional blocks, each consisting of two $\times3$ convolutional layers and two batch normalization layers. The number of channels in the five blocks are 64, 128, 256, 512, and 1024, respectively. The decoder contains four convolutional blocks, and the number of channels is set accordingly to the number of channels of the skip-connected features extracted by the encoder.
@@
  • For SiamCRNN, we applied pseudo-siamese ResNet-18 as encoder and four ConvLSTM layers (Shi et al., 2015) with $\times3$ convolutional kernel size and a hidden dimension of 128, as the decoder.
@@
Appendix E Details of post-processing methods
To further refine the raw predictions from DL models and enhance the quality of building damage maps, we apply three post-processing techniques: test-time augmentation (TTA), object-based majority voting, and model ensembling in Section 3.5. This section provides an overview of these methods and details the specific implementation used in our study.
@@ Appendix E Details of post-processing methods
  • Object-based majority voting aims to enforce spatial consistency by considering entire building instances rather than making independent pixel-wise predictions. Instead of classifying each pixel separately, the final label is determined based on a majority vote across the entire building object, leading to more coherent and reliable results. We follow the setup from (Zheng et al., 2021), where each building is treated as an independent object, and a weighted majority voting scheme is applied. The weighting is determined by the inverse of class proportions in the training set, meaning that underrepresented damage categories (e.g., destroyed buildings) have higher voting weights compared to dominant classes (e.g., intact buildings).
@@
Appendix F Details of unsupervised domain adaptation methods
@@
Appendix G Details of semi-supervised learning methods
@@
Final performance is evaluated on the target disaster event using mIoU as the primary evaluation metric.
Appendix H Unsupervised multimodal change detection: methods and evaluation protocol
Figure 13: Illustration of benchmark datasets commonly used in unsupervised multimodal change detection. The order from left to right is the corresponding order in the Table 13. In reference maps, white color indicates changed pixels, and black color indicates unchanged pixels. Shuguang, River, and Italy datasets are sourced from (Zhang et al., 2016), while the Gloucester dataset is from (Prendes et al., 2015).
H.1 Evaluated UMCD methods
@@
  • Random guessing assumes each pixel has an equal probability (50%) of being classified as changed or unchanged. This baseline provides a reference performance floor for evaluating the relative effectiveness of other UMCD methods.
@@
H.2 A more practical evaluation protocol on Bright
The training and evaluation protocols commonly adopted in the UMCD literature are often limited to individual image pairs (shown in Figure 13), using the same scene for representation learning, hyperparameter tuning, and evaluation. While this is understandable due to the lack of large-scale public datasets in this field, such a setup fails to reflect real-world use cases and often leads to overfitting and overestimated performance. To address this, we introduce a standardized and practical evaluation protocol using the Bright dataset. Specifically, we use the validation set from the standard ML split in Section 1.4 as the training set for UMCD methods, including any hyperparameter or threshold tuning. These models are then evaluated on the combined training and test sets, which are strictly held out during adaptation. This avoids overlap between tuning and evaluation scenes and better simulates the real-world deployment setting, where a model is expected to generalize to new, unseen data.
For evaluation, it is important to note that most of these methods were originally developed for land cover change detection, where changes correspond to transitions between semantic categories such as vegetation, water, or built-up areas. Applying them directly to disaster damage detection poses certain challenges. In particular, “Damaged” buildings do not always exhibit strong spatial or spectral signals, making them hard to distinguish in an unsupervised setting. However, “Destroyed” buildings often result in a complete change of land cover appearance (e.g., collapsed structures, debris, or ground exposure), which aligns better with the assumptions of existing UMCD methods. Therefore, in our evaluation, we treat buildings labeled as “Destroyed” in Bright as the positive (changed) class, and all other regions (including intact, damaged, and background) as the unchanged class. This ensures fair adaptation to disaster scenarios while respecting the original design of these methods.
Appendix I Details of unsupervised multimodal image matching methods
@@
In contrast, area-based methods, including FLSS (Ye et al., 2017) and HOPC (Ye et al., 2019), begin by identifying keypoints only in the optical image. These points are then matched by sliding a template window across a local search region in the SAR image, relying on local patch similarity. In our experiments, we used a template size of $\times120$ pixels, and a search window of $\times200$ pixels for both methods. This category is also implemented using open-source code [^13] https://github.com/yeyuanxin110/CFOG.
@@
where $(\hat{x}_{i}^{\text{sar}},\hat{y}_{i}^{\text{sar}})=\mathcal{T}(x_{i}^{\text{sar}},y_{i}^{\text{sar}})$ are transformed SAR coordinates.
@@
HC: conceptualization (lead), data curation (lead), funding acquisition (equal), methodology (lead), project administration (lead), investigation (lead), software (lead), visualization (lead), writing – original draft preparation (lead), writing – review and editing (equal). JS: conceptualization (support), data curation (equal), funding acquisition (support), methodology (support), investigation (support), software (support), writing – original draft preparation (equal), visualization (support). OD: conceptualization (support), data curation (equal), software (support), visualization (equal), writing – review and editing (support). CBB: conceptualization (support), data curation (support), writing – review and editing (lead). WX: data curation (support), project administration (support), writing – review and editing (equal). JW: data curation (support), writing – review and editing (support). XS: funding acquisition (support), writing – review and editing (support). YW: writing – review and editing (support). JX: conceptualization (support), data curation (support), supervision (support), writing – review and editing (support). CL [^14] Cuiling Lan did not participate in any activities related to the acquisition, processing, utilization, or distribution of the datasets. Her access to the data was solely limited to the information presented in the current paper, similar to that of the readers. : conceptualization (support), supervision (support), writing – linguistic refinement (equal). KS: conceptualization (equal), resources (support), supervision (equal), writing – review and editing (equal). NY: conceptualization (equal), funding acquisition (lead), resources (lead), supervision (lead), writing – review and editing (equal).